A large procession of people is moving along a road of width 10 m . There is a railway track across the road. The number of people present per square meter of road is 6 on an average . The average speed at which the procession is moving is 0.4 m/s . Find no of people crossing railway track per second

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
          Average area covered per second = 0.4 * 10 = 4 m^2 /s.
No. of people crossing the railway track per second = 4 * 6 = 24.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)